OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. Type your answers in the blanks. Show your work. |4x+3|=9+2x

OpenStudy (anonymous):

i got x=3 or x=-2 but she said my work was wrong again @KamiBug

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

So, we have two equations that we solve individually. 4x+3=9+2x 4x+3=−9−2x 4x+3=9+2x Subtract 3 from both sides. 4x=6+2x Subtract 2x from both. 2x=6 Divide both sides by 2. x=62=3 then i did the same for the second equation and got my answer but she said it was wrong

OpenStudy (anonymous):

@KamiBug @biby what did i do wrong she said my work for the second equation was wrong

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

what was your 2nd answer?

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

@Lakeekee_

OpenStudy (bibby):

I got that too, weird

OpenStudy (anonymous):

what did you do to get that answer

Directrix (directrix):

>> x=62=3 is not true. Look at it. If your teacher is looking at your work, this won't go over well. It is supposed to be 6/2. The work appears to be generated by some online site. I'm not saying it is. But, I bet the teacher thinks that and is looking at every little thing about the solution. @Tazmaniadevil